OpenStudy (anonymous):

Use Newton's method to find the absolute maximum value of the function f(x) = 3x sin x, 0 ≤ x ≤ π correct to six decimal places.

OpenStudy (anonymous):

@jlhurwit75 use the formula for newton method X,n+1 = X,n - f(x)/f'(x), start with n=0 so that X,1= X,0 - f(X0)/f'(x0) say X,o=3, then also try X,o =1 you might get a max of +pi, etc

OpenStudy (anonymous):

note this is an iterative process and do them successively for the approximation of roots

OpenStudy (anonymous):

yea i know that, I don't have a calculator with me and when I try to do it with mathmatica it doesnt work for this specific problem, I dont know the program well enough to know what I'm doing wrong